OpenStudy (tinash):

A mathematician states that her children’s ages are all prime numbers that multiply together to give 7,429. She also says that two of her children are teenagers. What are her children s age?

OpenStudy (freckles):

so we have \[a \cdot b \cdot c=7429\]

OpenStudy (freckles):

we know prime numbers are odd for the most part

OpenStudy (freckles):

and we know 7429 is not even so none of the primes are 2

OpenStudy (freckles):

two of her children are teenagers so they could be 13,17, or 19

OpenStudy (freckles):

I would probably try to write the prime factorization of 7429

OpenStudy (freckles):

can you do that?

OpenStudy (tinash):

wait what are prime numbers? :$

OpenStudy (freckles):

prime numbers are integers that are only divisible by themselves and 1.

OpenStudy (freckles):

the key here is just to write 7429 in prime factorization that is all there is to the problem

OpenStudy (freckles):

like for example 25=5*5 or 5^2 31 is prime 25 was not umm... 2 is the only even prime

OpenStudy (tinash):

I dont even know what factorization means >.< I missed few days of school.

OpenStudy (campbell_st):

one tip... given the last digit is 9... I'd divide the original number by 17

OpenStudy (freckles):

example: the prime factorization of 12=2*2*3

OpenStudy (tinash):

thanks <3

OpenStudy (campbell_st):

then see if the other 2 primes in the teens divides evenly

OpenStudy (tinash):

I don't know , I will try to figure it out thanks though

OpenStudy (freckles):

well the key to help was the teenager part

OpenStudy (tinash):

yeah , teens

OpenStudy (freckles):

"so they could be 13,17, or 19" exactly two of these numbers will be factors

OpenStudy (freckles):

then you have to find the other factor by divide 7429 by the product of two teenagers

OpenStudy (freckles):

what is 7429 divided by 13 ?

OpenStudy (tinash):

2476.3

OpenStudy (freckles):

you should get some kind of remainder that isn't 0 which would mean 13 is not factor of 7429 which means the other two number 17 and 19 are factors of 7429

OpenStudy (freckles):

so you only have to find one last factor

OpenStudy (freckles):

\[\frac{7429}{17 \cdot 19}=?\]

OpenStudy (tinash):

So far I have done , 17/7429 = 437 /23 = 19

OpenStudy (tinash):

that leaves me with 1 , if i divide it within 19 itself

OpenStudy (freckles):

is this what you meant to say: \[\frac{7429}{17 \cdot 19}=\frac{7429}{323}=23\]

OpenStudy (freckles):

which means 17*19*23=7429

OpenStudy (tinash):

kinda

OpenStudy (freckles):

oh I think I get what you are saying but those fractions aren't equal

OpenStudy (freckles):

\[\frac{7429}{17}=437 \\ \frac{437}{23}=19\]

OpenStudy (freckles):

first fraction gives 7429=17(437) then second fraction says 437=23(19)

OpenStudy (freckles):

so replace 437 in 7429=17(437) with 23(19) so 7429=17(23*19)

OpenStudy (freckles):

and you did it girl good job!

OpenStudy (tinash):

Oh my gosh thank you so much <3

OpenStudy (freckles):

and 17,19,23 are all prime because they are only divisible by themselves or 1

OpenStudy (tinash):

Can I get this straight first XD I wanna be completely sure 17/7421 = 19 divide by 23 or what? Oh my gosh I am confusing myself